OpenStudy (anonymous):

pls whoever gives me correct answer first gets medals plzz I need helllp Solve -4x - 5 > -9 and describe the graph of the solution. x > 1; closed circle on 1, shading to the right x < 1; closed circle on 1, shading to the left x < 1; open circle on 1, shading to the left x > 1; open circle on 1, shading to the right

OpenStudy (anonymous):

any 1

OpenStudy (anonymous):

@david1990 this one

OpenStudy (ivettef365):

first add 5 to both sides

OpenStudy (ivettef365):

what do you get ?

OpenStudy (anonymous):

-4x>-4

OpenStudy (ivettef365):

it is -4x - 5 + 5 > -9 + 5

OpenStudy (anonymous):

its either c or d

OpenStudy (ivettef365):

good, now divide by -4, since you are dividing by a negative number you need to switch the sign from > to <

OpenStudy (anonymous):

so its c?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

I have a couple more if you don't mind could you help me?

OpenStudy (ivettef365):

sure

OpenStudy (anonymous):

..

OpenStudy (ivettef365):

which question ?

OpenStudy (anonymous):

all

OpenStudy (ivettef365):

ok the first one you need to find the value of x on all the answer choices, I got for answer the last one which is 37 <= 8x -3 <= 61 +3 +3 +3 ------------------- 40 <= 8x <= 64 -- -- --- 8 8 8 5<= x <= 8 so x>= 5 and x<= 8 that is the graph

OpenStudy (ivettef365):

The next problem is more easy because you don't need to solve for x, is already given. remember open dot means (< or >), closed dot means (<= or >=) which one you think is the answer

OpenStudy (anonymous):

5x + 7 greater than or equal to 2 and 10x less than or greater to 30 6x + 7 greater than or equal to 25 or 2x - 8 < -4 4x + 10 less than or greater to 6 or 5x - 6 greater than or equal to 14 37 less than or greater to 8x - 3 less than or greater to 61

OpenStudy (anonymous):

these are the choices for number 1

OpenStudy (ivettef365):

already did one, look at the one I did

OpenStudy (anonymous):

..............

OpenStudy (ivettef365):

ok the first one you need to find the value of x on all the answer choices, I got for answer the last one which is 37 <= 8x -3 <= 61

OpenStudy (anonymous):

wb the other 2

OpenStudy (ivettef365):

The next problem is more easy because you don't need to solve for x, is already given. remember open dot means (< or >), closed dot means (<= or >=) which one you think is the answer

OpenStudy (anonymous):

D

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):



OpenStudy (anonymous):

wb the 3rd one

OpenStudy (anonymous):